# When Cars meet Drones: Hyperbolic Federated Learning for Source-Free Domain Adaptation in Adverse Weather

### v1

@@ Abstract
**footnotetext: These authors contributed equally to this work.
@@ 1 Introduction
The field of autonomous driving has evolved significantly beyond the traditional concept of cars navigating city streets. In the contemporary landscape, autonomous navigation encompasses various perspectives, including drones and robots, serving a wide range of applications like autonomous delivery and surveillance. The demand for intelligent systems that can seamlessly generalize across these varying viewpoints is quickly increasing, to enhance safety, efficiency, and adaptability. A key challenge lies in achieving this goal while preserving confidentiality and data security. The importance of robust privacy measures in developing autonomous systems cannot be overstated, as these systems often rely on substantial amounts of data to function effectively. The traditional method of gathering and analyzing data in a centralized manner raises relevant security concerns. Federated learning allows tackling this issue by introducing a cutting-edge paradigm that enables the collaborative training of machine learning models without sharing the private data of the participants (i.e., clients). In this paper, we introduce a novel federated learning approach tailored to the specific challenges encountered in driving scenarios, where autonomous agents have different viewpoints, specifically the street-view vision for cars and the aerial for UAVs. Multi-viewpoint paradigms have been proven beneficial in various tasks among which localization and segmentation [35, 31, 18], where jointly training the model with images from different viewpoints improves the objective task. Consequently, we enhance our framework by integrating aerial data. Moreover, we exploit this strategy in combination with the recent federated learning paradigm proposed in [37] that exploits a supervised server pretraining to enable unsupervised training on the client side. Additionally, our approach not only enables knowledge sharing between clients with different identities - either cars or drones - but also dynamically adapts to ever-changing weather conditions, aligning with the real-world challenges in this dynamic field. In this scenario (shown in Fig. 1), several key challenges arise:
  1. 1.
    Clients operating without supervision tend to become overly confident in their predictions, causing the model to diverge.
  2. 2.
    Clients exhibit differing distributions, and varying weather conditions can significantly impact performance (i.e., clients are non-iid, heterogeneous).
  3. 3.
    Clients are unbalanced, reflecting real-world scenarios where certain types of agents may be less common (e.g., fewer drones compared to cars).
Figure 1: Overview of the considered setting: after a supervised pretraining on synthetic data, a heterogeneous set of unsupervised clients collaboratively optimize their model in a federated fashion. Notably, adapting the model also to varying weather conditions.
@@
  • We introduced an unsupervised federated learning system for autonomous cars that can benefit from aerial view integration and seamlessly switch between the two perspectives.
  • We introduced HyperFLAW(Hyperbolic Federated Learning in Adverse Weather), which addresses domain shifts in diverse weather conditions using weather-aware batch normalization layers and integrates the first semantic segmentation dataset with adverse weather imagery for aerial vehicles.
  • We exploited prototype-based learning in the hyperbolic space to ensure consistent training across the clients in this very challenging setting.
@@ 2 Related Work
Domain Adaptation (DA) primarily focuses on adapting a model from a source to a target domain, typically in scenarios where the target data lacks labels, referred to as Unsupervised Domain Adaptation (UDA) [36, 43]. Historically, DA methods aimed to bridge the gap by quantifying domain divergence [45, 33]. Adversarial learning has also been widely used [44, 24, 30]. In the contemporary landscape, advanced approaches [54, 15, 28] leverage self-learning techniques. Several other approaches have been used to align the domains such as image-to-image translation [47, 1, 6], pseudo-label prototypes [51, 23, 50], and confidence thresholding [28, 54]. Within the objective of adapting to adverse weather conditions, Bruggemann et al. [3] propose a solution that requires sunny daytime references. To tackle nighttime images, Xiao et al. [46] employed a data augmentation method leveraging multiple modalities. Finally, Source Free Domain Adaptation (SFDA) presents a distinct advantage in addressing the performance gap by allowing adaptation without access to labeled source data [19].
Federated Learning has captured the interest of the research community due to its practical applicability in various real-world scenarios and its potential to manage sensitive data [27]. In literature, it has been applied to different vision tasks [38], initially with a large emphasis on classification [21, 20], but more recently also focusing on other tasks including semantic segmentation for Autonomous Driving [5, 11, 37]. Most existing works in FL assume labeled data on remote clients [5], which is an unrealistic assumption, especially in the driving scenario. Notably, Shenaj et al. [37] introduce the Federated source Free Domain Adaptation (FFreeDA) setting, where only the server accesses a labeled source dataset. We address this challenge by incorporating the management of diverse viewpoints, accommodating various client types, and dealing with the influence of adverse weather conditions.
@@ 3 Problem Formulation
Most works in Federated Learning operate under the assumption of labeled data being available on remote clients [5]. Providing semantic segmentation ground truth is extremely time-consuming and costly, thus it is quite unrealistic to expect labeled data on real-world clients, especially during the deployment phase. Therefore, our objective is to develop a model that can seamlessly adapt from labeled synthetic data to unlabeled real-world images in a distributed paradigm. For this reason, we chose to deploy our setting on top of the pre-existing FFreeDA scenario [37]. Moreover, we discard the assumption that the clients are homogeneous in nature and consider a setting where different autonomous agents work cooperatively in fluctuating atmospheric conditions (i.e., clear sky, night, rain and fog). In the following, we will first formalize the problem and then move to a detailed description of the methodologies proposed to tackle it in Sec. 4.
@@
Multi-viewpoint FFreeDA in Adverse Weather. Aiming to assist agents lacking large training datasets or capabilities, such as drones, we consider a setting that enables scene understanding for unsupervised autonomous navigation robust to multiple viewpoints. First of all, we consider $\mathcal{D}^{S}=\mathcal{D}^{S}_{car}\cup\mathcal{D}^{S}_{drone}$. Accounting for adverse weather as a heterogeneity factor, our dataset $\mathcal{D}^{S}=\{(\mathbf{x}^{S},\mathbf{y}^{S},w^{S})\}^{N^{S}}$ also includes $w^{S}$, denoting the atmospheric condition for each frame, together with the color images $\mathbf{x}^{S}$ and the ground truth segmentation maps $\mathbf{y}^{S}$. Similarly, the client can accommodate either car or drone agents: $\mathcal{D}^{T}=\mathcal{D}^{T}_{car}\cup\mathcal{D}^{T}_{drone}$, although clients datasets are still local and disjoint. As in the previous setting, each client $k\in\mathcal{K}$ have its own set of unsupervised images both in terms of semantics and weather: $\mathcal{D}^{T}_{k}=\{(\mathbf{x}^{T}_{k})\}$. Notably, the number of samples of the drones’ target dataset ($N^{T}_{drone}$) is typically lower than in the target datasets for cars ($N^{T}_{car}$). Furthermore, the set of semantic classes could differ between car and aerial vehicles datasets (in our case $C_{drone}\subset C_{car}=C^{T}=C^{S}$). The identity of each client is fixed, meaning it is designated as either a car or a drone and does not change during training, however, there is no such limitation on the weather conditions, which can change dynamically during the training process. The goal is to minimize the objective function of Eq. 1, where the target data distribution holds into its heterogeneity different viewpoints and weathers. In Fig. 2 we present our approach, which will be detailed in Sec. 4.
@@ 4 Training Procedure
Pretraining. We train the initial model $M^{0}_{G}$ on a large-scale source synthetic dataset $\mathcal{D}^{S}$, taking advantage of the server’s computing capabilities. By sampling batches of cross-agent we maximize the standard cross entropy loss $\mathcal{L}(y^{S},M_{G}^{0}(x^{S}))$. Moreover, an additional component $M_{W}$ consisting of 3 convolutional layers is trained to recognize weather conditions by optimizing $\mathcal{L}(w^{S},M_{W}(x^{S}))$. This classifier is a lightweight module intended to provide weather information for clients and it is not trained after this phase. Differently from [37], we do not assume to have information on the clients’ distribution (i.e., in [37] the average Fourier domain coefficients from $\mathcal{D}^{T}$ are used, requiring priors on clients’ data).
Adaptation. At each round $r$, we consider that, among all the set of clients $\mathcal{K}$, only a subset denoted as $\mathcal{K}^{r}$ participate in the training. This assumption is common in FL where not all the clients might be concurrently active and reflects well on our scenario (e.g. a car is parked). Then, the server transmits the global model from the previous round $M_{G}^{r-1}$. Note that the weather classifier $M_{W}$ is only transmitted once as it is fixed. The training for the active clients will be performed separately on $M_{k}^{r},k\in\mathcal{K}^{r}$. Furthermore, a prototype representation $\mathbf{p_{c,k}^{r}}$ for each class (i.e., the centroids of the class samples in feature space) is also built and transmitted to the server together with the manifold curvature $\gamma_{k}$ (refer to Sec. 4.3).
@@
Finally, the models are aggregated at the server side with the approach of Sec. 4.4 obtaining the global model ${M}_{G}^{r}$ that is used as input for the next step. Prototypical representations are also aggregated at the server side obtaining a set of global class prototypes $\mathbf{p^{r}_{G,c}}$.
@@ 4.1 Client Self-Training (ST)
Self-training involves iteratively enhancing a model’s performance by using its own predictions. In FL, the clients can rely on the server model as the teacher, as ideally, the server aggregates the global knowledge from the targets. In our framework, we update the teacher model after the round is over. This implies that the global model received by the clients at the beginning of each round $r$, $M_{G}^{r-1}$, serves also as the current teacher. This approach offers the advantage of eliminating the necessity to transmit two models per round as in [37] where also a model from $R_{u}$ steps before was used, thus requiring only a single model transmission and avoiding the extra hyperparameter $R_{u}$. The local model of the $k-th$ client is trained using the weak supervision of the pseudo-labels produced by the teacher model $M_{G}^{r-1}$ and its local model prediction $M^{r}_{k}$ by optimizing the self-training loss obtained via cross-entropy as $\mathcal{L}_{st}(M_{G}^{r-1}(x_{k}^{T}),M^{r}_{k}(x_{k}^{T}))$. Moreover, to achieve a training that is consistent across the heterogeneous clients and a model able to generalize to different types of agents and weather conditions the self-training loss is used in combination with the prototype-based one, i.e.,
@@ 4.2 Weather Batch Normalization (BN)
Inspired by [22] where the update of the batch normalization layers is not synchronized within the global model, we personalize the training by considering a different set of batch-normalization (BN) parameters for each weather condition $w$, defined as $\{\theta^{BN}_{w}\}$. After the pretraining ($r=0$), all the weather-specific BNs inherit the global BN statistics $\theta^{BN}$, while in each subsequent round, the statistics of the BNs are averaged only over those of the same weather condition. Specifically, on the client side, after obtaining the prediction $\bar{w}=M_{W}(x^{T})$ from the classifier, the client batch-norm parameters are optimized using Eq. 2 only on the selected weather-BNs $\theta^{BN}_{\bar{w}}$. This allows us to enforce a more general feature extraction from the convolutional layers while exploiting the domain-specific nuances of road scenes under different weather conditions.
@@ 4.3 Prototypical Learning
At the beginning of each round $r$, a participating client $k$ initializes its local prototypes $\mathbf{p_{k,c}^{r}}$ using the previous round ones received from the server: $\mathbf{p_{k,c}^{r}}=\mathbf{p_{G,c}^{r-1}}$ Then, the prototypes are updated at training time using exponential smoothing with rate $\beta=0.85$ to stabilize their evolution as follows:
$\mathbf{p_{c}}=\beta\;\mathbf{p^{\prime}_{c}}+(1-\beta)\;\mathbf{p_{c}}$ (4)
where $\mathbf{p^{\prime}_{c}}$ is the prototype from the previous step. After each round, prototypes $\mathbf{p_{k,c}^{r}},\forall c\in\mathcal{C}$ computed by each client are sent back to the server for aggregation.
@@
$d_{\gamma}(\mathbf{x},\mathbf{y})=\left(\frac{2}{\sqrt{\gamma}}\right)\tanh^{-1}\left(\sqrt{\gamma}\|\mathbf{-x}\oplus_{\gamma}\mathbf{y}\|\right)$ (6)
@@
Prototypical Loss. Due to the distribution mismatch between source and target domains and the different viewpoints, feature vectors from these domains become misaligned, leading to inaccuracies in target representations and a subsequent decline in segmentation accuracy. To address this issue, we adopt a clustering objective in the latent space by leveraging prototypical representations:
$\mathcal{L}_{cl}=\frac{1}{|\hat{\mathcal{C}}|}\sum_{c\in\hat{\mathcal{C}}}\frac{1}{|\mathcal{F}_{c}|}\sum_{\mathbf{f_{c}\in\mathcal{F}_{c}}}d(\mathbf{f}_{c},\mathbf{p}_{c})$ (11)

### v2

@@ 1 Introduction
The field of autonomous driving has evolved significantly beyond the traditional concept of cars navigating city streets. In the contemporary landscape, autonomous navigation encompasses various perspectives, including drones and robots, serving a wide range of applications like autonomous delivery and surveillance. The demand for intelligent systems that can seamlessly generalize across these varying viewpoints is quickly increasing, to enhance safety, efficiency, and adaptability. A key challenge lies in achieving this goal while preserving confidentiality and data security. The importance of robust privacy measures in developing autonomous systems cannot be overstated, as these systems often rely on substantial amounts of data to function effectively. The traditional method of gathering and analyzing data in a centralized manner raises relevant security concerns. Federated learning allows tackling this issue by introducing a cutting-edge paradigm that enables the collaborative training of machine learning models without sharing the private data of the participants (i.e., clients).
Figure 1: Federated Learning for autonomous driving across diverse conditions: each client trains the segmentation model locally with its own data, which may be biased towards specific environmental conditions. E.g., a taxi operating predominantly at night will have most training samples in nighttime conditions. Despite the local biases, the approach aims to create a single, robust model capable of generalizing across diverse driving scenarios.
In this paper, we introduce a novel federated learning approach tailored to the specific challenges encountered in driving scenarios, where autonomous agents have different perspectives, specifically the street-view vision for cars and the aerial for UAVs. Multi-viewpoint paradigms have been proven beneficial in various tasks including localization and segmentation [32, 9, 16], where jointly training the model with images from different viewpoints improves the objective task. Consequently, we enhance our framework by integrating aerial data during the federated learning process. Importantly, while our approach leverages multi-viewpoint data to learn a more generalizable model, it does not require matching views at inference or training time for individual clients. Each client operates solely with its own perspective, yet benefiting from the collective knowledge gained through federated learning across diverse viewpoints.
Additionally, we dynamically adapt to ever-changing weather conditions, aligning with the real-world challenges in this dynamic field. We combine this strategy with the federated paradigm proposed in [34], which employs supervised server pretraining to enable unsupervised client-side training. In this scenario (shown in Fig. 1), several key challenges arise: 1 Clients operating without supervision tend to become overly confident in their predictions, causing the model to diverge; 2 Clients exhibit differing distributions, and varying weather conditions can significantly impact performance (i.e., clients are non-iid, heterogeneous); 3 Clients are unbalanced, reflecting real-world scenarios where certain types of agents may be less common (e.g., fewer drones compared to cars).
@@
  • We introduce an unsupervised federated system for learning a global model which can generalize between cities, weather and perspectives.
  • We provide FLYAWARE, the first semantic segmentation dataset with adverse weather imagery for aerial vehicles.
  • We propose HyperFLAW  [^1] The code and dataset are available at https://github.com/LTTM/HyperFLAW (Hyperbolic Federated Learning in Adverse Weather), which addresses domain shifts in diverse weather conditions using weather-aware batch normalization layers.
  • We exploit prototype-based learning in the hyperbolic space to ensure consistent training across the clients in this very challenging setting.
@@ 2 Related Work
Domain Adaptation (DA) primarily focuses on adapting a model from a source to a target domain, typically in scenarios where the target data lacks labels, referred to as Unsupervised Domain Adaptation (UDA) [33, 38]. Historically, DA methods aimed to bridge the gap by quantifying domain divergence [40, 30]. Adversarial learning has also been widely used [39, 22, 28]. In the contemporary landscape, advanced approaches [48, 14, 26] leverage self-learning techniques. Several other approaches have been used to align the domains such as image-to-image translation [44, 2, 7], pseudo-label prototypes [46, 21, 45], and confidence thresholding [26, 48]. Within the objective of adapting to adverse weather conditions, Bruggemann et al. [4] propose a solution that requires sunny daytime references. To tackle nighttime images, Xiao et al. [43] employed a data augmentation method leveraging multiple modalities. Finally, Source-Free Domain Adaptation presents a distinct advantage in addressing the performance gap by allowing adaptation without access to labeled source data [17].
Federated Learning has captured the interest of the research community due to its practical applicability in various real-world scenarios and its potential to manage sensitive data [25]. In literature, it has been applied to different vision tasks [35], initially with a large emphasis on classification [19, 18], but more recently also focusing on other tasks including semantic segmentation for Autonomous Driving [6, 11, 34]. Most existing works in FL assume labeled data on remote clients [6, 11], which is an unrealistic assumption, especially in the driving scenario. Notably, Shenaj et al. [34] introduce the Federated source Free Domain Adaptation (FFreeDA) setting, where only the server accesses a labeled source dataset. We address this challenge by incorporating the management of diverse viewpoints, accommodating various client types, and dealing with the influence of adverse weather conditions.
@@ 3 Problem Formulation
Most works in Federated Learning operate under the assumption of labeled data being available on remote clients [6, 11]. Providing semantic segmentation ground truth is extremely time-consuming and costly, thus it is quite unrealistic to expect labeled data on real-world clients, especially during the deployment phase. Therefore, our objective is to develop a model that can seamlessly adapt from labeled synthetic data to unlabeled real-world images in a distributed paradigm. For this reason, we chose to deploy our setting on top of the pre-existing FFreeDA scenario [34]. Moreover, we discard the assumption that the clients are homogeneous in nature and consider a setting where different autonomous agents work cooperatively in fluctuating atmospheric conditions (i.e., clear sky, night, rain and fog). In the following, we will first formalize the problem and then move to a detailed description of the methodologies proposed to tackle it in Sec. 4.
@@
Multi-viewpoint FFreeDA in Adverse Weather. Aiming to assist agents lacking large training datasets or capabilities, such as drones, we consider a setting that enables scene understanding for unsupervised autonomous navigation robust to multiple viewpoints. First of all, we consider $\mathcal{D}^{S}=\mathcal{D}^{S}_{car}\cup\mathcal{D}^{S}_{drone}$. Accounting for adverse weather as a heterogeneity factor, the source data are sampled as $\{(\mathbf{x}^{S},\mathbf{y}^{S},w^{S})\}^{N^{S}}$ which includes $w^{S}$, denoting the atmospheric condition for each frame, together with the color images $\mathbf{x}^{S}$ and the ground truth segmentation maps $\mathbf{y}^{S}$. Similarly, the client can accommodate either car or drone agents: $\mathcal{D}^{T}=\mathcal{D}^{T}_{car}\cup\mathcal{D}^{T}_{drone}$, although client datasets are still local and disjoint. Please note that within the target set, the drone and car data are considered to have been acquired separately, i.e., they belong to distinct domains. This assumption reflects the practical constraints of real-world scenarios, where obtaining matching viewpoints between aerial and terrestrial imagery is typically infeasible.
As in the previous setting, each client $k\in\mathcal{K}$ have its own set of unsupervised images both in terms of semantics and weather. Notably, the number of samples of the drones’ target dataset ($N^{T}_{drone}$) is typically lower than in the target datasets for cars ($N^{T}_{car}$). Furthermore, the set of semantic classes could differ between car and aerial vehicles datasets (in our case $C_{drone}\subset C_{car}=C^{T}=C^{S}$). The identity of each client is fixed, meaning it is designated as either a car or a drone and does not change during training, however, there is no such limitation on the weather conditions, which can change dynamically during the training process. The goal is to minimize the objective function of Eq. 1, where the target data distribution holds into its heterogeneity different viewpoints and weathers. In Fig. 2 we present our approach, which will be detailed in Sec. 4.
@@ 4 Training Procedure
Pretraining. We train the initial model $M^{0}_{G}$ on a source synthetic dataset $\mathcal{D}^{S}$, taking advantage of the server’s computing capabilities. By sampling batches of cross-agent we maximize the standard cross entropy loss $\mathcal{L}(y^{S},M_{G}^{0}(x^{S}))$. Moreover, an additional component $M_{W}$ consisting of 3 convolutional layers is trained to recognize weather conditions by optimizing $\mathcal{L}(w^{S},M_{W}(x^{S}))$. This classifier is a lightweight module intended to provide weather information for clients and it is not trained after this phase. Differently from [34], we do not assume to have information on the clients’ distribution (i.e., in [34] the average Fourier domain coefficients from $\mathcal{D}^{T}$ are used, requiring priors on clients’ data).
Adaptation. At each round $r$, we consider that, among all the set of clients $\mathcal{K}$, only a subset denoted as $\mathcal{K}^{r}$ participate in the training. This assumption is common in FL where not all the clients might be concurrently active and reflects well on our scenario (e.g. a car is parked). Then, the server transmits the global model from the previous round $M_{G}^{r-1}$. Note that the weather classifier $M_{W}$ is only transmitted once as it is fixed. The training for the active clients will be performed separately on $M_{k}^{r},k\in\mathcal{K}^{r}$. Client-side training employs self-training (Sec. 4.1), weather-batch normalization to align feature styles across climate conditions (Sec. 4.2), and, finally, prototypical learning for feature space generalization (Sec. 4.3). Each client generates class prototypes $\mathbf{p_{c,k}^{r}}$, i.e., centroids of class samples in feature space, and computes the manifold curvature $\gamma_{k}$, both of which are transmitted to the server. The server aggregates the client models as described in Sec. 4.4, producing the global model ${M}_{G}^{r}$ for the next round $r+1$. Global class prototypes $\mathbf{p^{r}_{G,c}}$ are also aggregated from the client prototypes.
@@ 4.1 Client Self-Training (ST)
Self-training involves iteratively enhancing a model’s performance by using its own predictions. In FL, the clients can rely on the server model as the teacher, as ideally, the server aggregates the global knowledge from the targets. In our framework, we update the teacher model after the round is over. This implies that the global model received by the clients at the beginning of each round $r$, $M_{G}^{r-1}$, serves also as the current teacher. This approach offers the advantage of eliminating the necessity to transmit two models per round as in [34] where also a model from $R_{u}$ steps before was used, thus requiring only a single model transmission and avoiding the extra hyperparameter $R_{u}$. The local model of the $k-th$ client is trained using the weak supervision of the pseudo-labels produced by the teacher model $M_{G}^{r-1}$ and its local model prediction $M^{r}_{k}$ by optimizing the self-training loss obtained via cross-entropy as $\mathcal{L}_{st}(M_{G}^{r-1}(x_{k}^{T}),M^{r}_{k}(x_{k}^{T}))$. Moreover, to achieve a training that is consistent across the heterogeneous clients and weather conditions, the self-training loss is used in combination with the prototype-based one, i.e.,
@@ 4.2 Weather Batch Normalization (BN)
Recent methods [20, 1] have explored unsynchronized batch normalization layers to tailor local clusters. Although offering local personalization, these approaches may be suboptimal in driving scenarios, especially when considering real-world applications. Imagine a situation where each client predominantly drives in specific weather conditions, resulting in training samples primarily drawn from a single weather distribution. Ideally, the model should perform equally well across all weather conditions, not just those frequently encountered by individual clients. To address this challenge, we propose a novel strategy: rather than customizing for each client, we adapt the training process for each distinct weather condition. This is accomplished by utilizing a separate set of batch-normalization parameters ${\theta^{BN}_{w}}$ for each weather condition $w$ (as illustrated in Fig. 3).
Figure 3: Comparison of Weather-Batch Normalization with BN personalization methods [20, 1].
After the pretraining ($r=0$), all the weather-specific BNs inherit the global BN statistics $\theta^{BN}$, while in each subsequent round, the statistics of the BNs are averaged only over those of the same weather condition. Specifically, on the client side, after obtaining the prediction $\bar{w}=M_{W}(x^{T})$ from the classifier, the client batch-norm parameters are optimized using Eq. 2 only on the selected weather-BNs $\theta^{BN}_{\bar{w}}$. This allows us to enforce a more general feature extraction from the convolutional layers while exploiting the domain-specific nuances of road scenes under different weather conditions.
@@ 4.3 Prototypical Learning
At the beginning of each round $r$, a participating client $k$ initializes its local prototypes $\mathbf{p_{k,c}^{r,0}}$ using the previous round ones received from the server: $\mathbf{p_{k,c}^{r,0}}=\mathbf{p_{G,c}^{r-1}}$. Then, at each training step $s$ the class prototypes at client side $\mathbf{p_{k,c}^{r,s}}$ are updated using exponential smoothing with rate $\beta=0.85$ to stabilize their evolution as follows:
$\mathbf{p_{k,c}^{r,s}}=\beta\;\mathbf{p_{k,c}^{r,s-1}}+(1-\beta)\;\mathbf{p_{k,c}^{r,s}}$ (4)
where $\mathbf{p_{k,c}^{r,s-1}}$ is the prototype from the previous training step. After each round, prototypes $\mathbf{p_{k,c}^{r}},\forall c\in\mathcal{C}$ computed by each client are sent back to the server for aggregation.
@@
$d_{\gamma}(\mathbf{x},\mathbf{y})=2\gamma^{-1/2}\tanh^{-1}\left(\sqrt{\gamma}\|\mathbf{-x}\oplus_{\gamma}\mathbf{y}\|\right)$ (6)
@@
Prototypical Loss. Due to the distribution mismatch between source and target domains and the different viewpoints, feature vectors from these domains become misaligned, leading to inaccuracies in target representations and a subsequent decline in segmentation accuracy. To address this issue, we adopt a clustering objective in the latent space by leveraging prototypical representations that exploits the distance between each extracted feature vector and the prototype available at client side for the corresponding class:
$\mathcal{L}_{cl}=\frac{1}{|\hat{\mathcal{C}}|}\sum_{c\in\hat{\mathcal{C}}}\frac{1}{|\mathcal{F}_{c}|}\sum_{\mathbf{f_{c}\in\mathcal{F}_{c}}}d(\mathbf{f}_{c},\mathbf{p}_{k,c}^{r})$ (11)